The ratio between the neutrons in C and Si with respect to atomic masses 12 and 28 is

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Pallavi,

C atomic mass is 12...implies no of neutrons = 6

Si atomic mass is 28...implies no of neutrons = 14

ratio of neutrons=6/14=3/7